x-6y=-7
9x-8y=-17

Answer by mananth(16946)

1 x -6 y = -7 .............1
9 x -8 y = -17 .............2
Eliminate y
multiply (1)by 8
Multiply (2) by -6
8 x -48 y = -56
-54 x 48 y = 102
Add the two equations
-46 x = 46
/ -46
x = -1
plug value of x in (1)
1 x + -6 y = -7
-1 + -6 y = -7
-6 y = -7 + 1
-6 y = -6
y = 1